Electromagnetic radiation of frequency 9.0 x 10^14Hz is incident on a clean metal surface. The work function of the metal is 5.0 x 10^-19J. Determine the maximum kinetic energy of photoelectrons released from the metal surface


1
Expert's answer
2021-08-20T18:04:43-0400

The photoelectric equation is

"\frac{hc}{λ} = ϕ+E_K"

EK=Kinetic energy

h=plank's constant

c=speed of light

λ=wavelength

ϕ=work function

"E_K = \frac{hc}{λ} -ϕ

λ = \frac{c}{frequency} \\

= \frac{3 \times 10^8}{9.0 \times 10^{14}} \\

= 3.33 \times 10^{-7} \\

h=6.626 \times 10^{-34} \;Js \\

c= 3 \times 10^8 \;m/s \\

ϕ = 5.0 \times 10^{-19} \;J \\

E_K = \frac{6.626 \times 10^{-34} \times 3 \times 10^8}{3.33 \times 10^{-7}} -5.0 \times 10^{-19} \\

= 5.969 \times 10^{-19} -5.0 \times 10^{-19} \\ 

= 0.969 \times 10^{-19} \;J \\

≈ 1 \times 10^{-19} \;J"